How much pure acid should be mixed with 5 gallons of a 70% acid solution in order to get a 90% acid solution?
Answer by mananth(16946)

--------percent ---------------- quantity
Acid solution 70 ---------------- 5 gallons
Pure acid 100 ---------------- x gallons
Mixture 90 ---------------- 5 + x gallons

70 * 5 + 100 x = 90 ( 5 + x )
350 + 100 x = 450 + 90 x
100 x + -90 x = 450 - -350
10 x = 100
/ 10
x = 10
10 gallons of Pure acid